OpenStudy (anonymous):

one leg of a right triangle measures 15 m and the hypotenuse measures 17 m. what is the measure of the other leg. A. 22.67 B. 8 C. 32 m D. 5.66 m

OpenStudy (anonymous):

Do you know what the Pythagorean Theorem is?

OpenStudy (anonymous):

... if not, its is A^2 +B^2 = C^2..... so to figure this problem out, just plug the numbers in

OpenStudy (anonymous):

okay well my bad for posting this but i got the answer but can u make sure? i got 8

OpenStudy (anonymous):

well B

OpenStudy (anonymous):

okay, its fine. Let me check.

OpenStudy (anonymous):

x^2 + 15^2 = 17^2 x = 8 So, B is correct i got this

OpenStudy (anonymous):

it is not B... theres something wrong with your math. try again

OpenStudy (anonymous):

whats 15^2 and 17^2?

OpenStudy (anonymous):

no wait is it A? 22.67 m?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

when you do 15^2 it is 225, 17^2 is 285... when you add those two it is 514. then you do the sqare root of 514 which is 22.67

OpenStudy (anonymous):

15^2 = 225 and 17^2 = 289 289-255 = 34 So the other leg is of length \[\sqrt{34}\]

OpenStudy (anonymous):

oh ok, thank you @Jamie_elizabeh :)

OpenStudy (anonymous):

yourwelcome:)

OpenStudy (anonymous):

@caters .. you have to add the two numbers not subtract. After you add the numbers, then find the square root of the numbers

OpenStudy (anonymous):

\[\sqrt{34} = 5.8309...\]

OpenStudy (anonymous):

no, add 15^2 and 17^2 then find the square root. You subtracted the two numbers. thats where you went wrong

OpenStudy (anonymous):

@caters

OpenStudy (anonymous):

He is wanting the length of the other leg. You add when you want the hypotenuse and subtract when you want the length of one of the legs so that is why I subtracted 17^2 and 15^2

OpenStudy (anonymous):

\[A^2 + B^2 = C^2\] so \[C^2 - A^2 = B^2\] and \[C^2 - B^2 = A^2\]

OpenStudy (anonymous):

....its the same thing whether you want the other leg or the hypotenuse... or at least i though it was.. let me google it..

OpenStudy (anonymous):

http://www.mathopenref.com/pythagorastheorem.html you dont subtract:)

OpenStudy (anonymous):

By adding 17^2 and 15^2 you get the hypotenuse of a bigger triangle. He explicitly stated that the hypotenuse is 17 and one of the legs is 15 so you subtract 17^2 and 15^2 to get the other leg.

OpenStudy (anonymous):

okay. i get how that would work but ima stick with the original formula